# Decomposing Multivariate Information Rates in Networks of Random Processes

### v2

@@ IV.2 Effects of changes in the network topology
$Y_{n}=(0.8-c)X_{1,n-1}+(1.6-2c)X_{2,n-1}+U_{Y,n}$ (33)
$X_{1,n}=cY_{n-1}+U_{X_{1},n}$
$X_{2,n}=2cY_{n-1}+U_{X_{2},n}$
@@
where $U_{Y}$, $U_{X_{1}}$ and $U_{X_{2}}$ are uncorrelated Gaussian white noises with zero mean and unit variance; the uncorrelation between the inputs ($\mathbf{\Sigma}_{U}=\mathbf{I}$) denotes absence of instantaneous correlations. On the other hand, causal interactions are set at lag $k=1$, from $Y$ to both $X_{1}$ and $X_{2}$ with strength modulated directly by the parameter $c$, and from $X_{1}$ and $X_{2}$ to $Y$ with strength modulated inversely by the same parameter $c$. This setting allows for a progressive strengthening of the causal interactions directed from the target to the sources, and a progressive weakening of the causal interactions directed from the sources to the target, as $c$ increases in the range $[0-0.8]$. The simulation design is shown in Fig. 4a.
@@
The PIRD was performed as in the first simulation, using the approach presented in Sect. "Formulation for Gaussian processes". Remarkably, since in this simulation of processes without self-dependencies the spectral MIR functions are flat, the time- and frequency-domain PIRD are equivalent and reduce to applying the MMI criterion to the MIR. As regards the PID applied to the joint TE $I(Y_{n};X_{1,<n},X_{2,<n}|Y_{<n})$, it was computed exploiting the formalism linking information-theoretic measures with Granger causality measures derived from linear parametric regression models developed for joint Gaussian processes Barnett et al. (2009b); Barrett et al. (2010); Faes et al. (2017). Specifically, the MMI criterion Barrett (2015) was used to derive redundancy as the minimum information transferred from each individual source to the target; the latter was measured as half the value of the Granger causality (GC) estimated from the VAR processes Barnett et al. (2009b), and all GC terms were computed from the VAR parameters using sub-models Faes et al. (2017). The resulting redundant TE ($\mathcal{R}$) was then used to derive the unique ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$) and synergistic ($\mathcal{S}$) amounts of information transferred from $X$ to $Y$ according to PID rules.
The time domain behaviors of the PID terms are shown in Fig. 4b-f as a function of the parameters $c_{1},c_{2}$ (continuous black lines).
@@
The results evidence how the transition from the condition in which the target acts exclusively as a sink of information ($c=0$) to that in which it acts exclusively as a source of information ($c=0.8$) is thoroughly reflected by the PIRD but cannot be fully captured by the PID applied to the joint TE. Indeed, the latter cannot take time-lagged interactions directed from the target to the sources into account, and thus in the simulation where $c$ is increased it reflects only the decrease of the joint information transferred along the direction $X\rightarrow Y$ ($J$, Fig. 4d, solid line) and the consequent drop to zero of all PID terms ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$, $\mathcal{R}$, $\mathcal{S}$, Fig. 4b,c,e,f, solid lines). On the other hand, the PIRD is applied to the MIR shared between the sources and the target, a measure which is not directional as the TE and is thus sensible to the overall information flowing among the processes of the considered dynamic system. This is reflected by values of the MIR between $Y$ and $X_{1},X_{2}$ which remain high as $c$ increases ($J$, Fig. 4d, dashed line), as also happens for the unique information rate of $X_{2}$ ($\mathcal{U}_{2}$, Fig. 4c, dashed line), and by values of redundancy rate and synergy rate shifting from the prevalence of synergy when $c=0$ to the prevalence of redundancy when $c=0.8$. The latter behavior reflects the modifications of the topological structure induced in the dynamic network, with links undergoing a transition from a common child configuration to a common drive configuration.
@@
Figure 4: Changes in the network topology have a profound impact on the multivariate information shared by multiple random processes. a) Simulation design, where $Y$ is the target process and $\{X_{1},X_{2}\}$ is the group of sources; time-lagged interactions (solid black arrows) are set varying the parameter $c\in[0-0.8]$ to simulate a transition from purely unidirectional interactions directed from $X_{1},X_{2}$ to $Y$ when $c=0$ to purely unidirectional interactions directed from $Y$ to $X_{1},X_{2}$ when $c=0.8$. (b-e) Profiles of the time domain MIR shared between the $Y$ and $\{X_{1},X_{2}\}$ and the joint TE from $\{X_{1},X_{2}\}$ to $Y$ (measuring overall interactions, here indicated as $J$) and of the unique ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$), redundant ($\mathcal{R}$) and synergistic ($\mathcal{S}$) components of their decomposition measured using the PID applied to the TEs (continuous black lines) and the PIRD (dashed gray lines) as a function of the parameter $c$.
@@ IV.3 Frequency-specific coarse-grained PIRD
Figure 5: Coexistence of redundant and synergistic characters of interactions in different spectral bands elicited by frequency-specific PIRD. (a) Network structure, with $Y$ receiving from $X_{1}$, oscillating at 0.3 Hz, and $X_{3}$, oscillating at 0.1 Hz, and $X_{1}$ sending to $X_{2}$, oscillating at 0.3 Hz; all coupling coefficients are set to 1. (b) Spectral profiles of the pairwise MIR measures computed between the target $Y$ and each source $X_{i}$, $i=1,2,3$. (c,d) Time-domain values of the pairwise MIR and of the unique ($\mathcal{U}_{Y;X_{i}}$, $i=1,2,3$), redundant ($\mathcal{R}_{Y;X}$) and synergistic ($\mathcal{S}_{Y;X}$) PIRD components integrated along the whole frequency axis ($\left[0-f_{s}/2\right]$) (left bars), and withing the bands $B_{1}=\left[0.04-0.15\right]$ Hz (middle bars), $B_{2}=\left[0.15-0.4\right]$ Hz (right bars); The balance $\Delta_{Y;X}=\mathcal{R}_{Y;X}-\mathcal{S}_{Y;X}$ is also reported, indicating overall prevalence of redundancy and coexistence of redundancy and synergy in the two different bands $B_{1}$ and $B_{2}$.
@@
In the last simulation, we provide an illustrative example of the full-frequency and band-limited PIRD applied to a network of four nodes where different high-order behaviors emerge at different frequencies. We consider a VAR process simulated with $f_{s}=1$, where the lagged interactions between the target $Y$ and the three sources $X=\left[X_{1},X_{2},X_{3}\right]$ are set to induce a network topology with both common drive and common child structures (Fig. 5a). The VAR process is defined as:
$Y_{n}=X_{1,n-1}+X_{3,n-1}+U_{Y,n}$ (34)
$X_{1,n}=\sum_{k=1}^{2}a_{1,k}X_{1,n-k}+U_{X_{1},n}$
$X_{2,n}=\sum_{k=1}^{2}a_{2,k}X_{2,n-k}+X_{1,n-1}+U_{X_{2},n}$
$X_{3,n}=\sum_{k=1}^{2}a_{3,k}X_{3,n-k}+U_{X_{3},n}$
@@
where $U=\left[U_{Y},U_{X_{1}},U_{X_{2}},U_{X_{3}}\right]$ is a vector of four zero-mean independent Gaussian white noises with unit variance. The autonomous oscillations in the three source processes $X_{1}$, $X_{2}$, $X_{3}$ are obtained placing a pair of complex-conjugate poles, with modulus $\rho$ and phase $2\pi{f}$, in the complex plane representation of each process. Here, we set $\rho=0.8,f=0.3$ Hz for the processes $X_{1}$ and $X_{2}$, so that their autonomous dynamics are determined by the coefficients $a_{1,1}=a_{2,1}=-0.494;a_{1,2}=a_{2,2}=-0.64$; similarly, we set $\rho=0.9,f=0.1$ Hz for $X_{3}$, so that its autonomous dynamics are determined by the coefficients $a_{3,1}=1.456;a_{3,2}=-0.81$. Moreover, causal interactions are set from $X_{1}$ and $X_{3}$ to $Y$ and from $X_{1}$ to $X_{2}$ at lag $k=1$, with unitary strength.
@@
The PIRD was applied computing the spectral MIRs between each individual source and the target ($i_{Y;X_{i}}(f)$, $i=1,2,3$; Fig. 5b), as well as between groups of sources and the target (e.g., $i_{Y;X_{1},X_{2}}(f)$); the spectral redundancy function was then computed as in (21), and was exploited to retrieve the unique ($\mathcal{U}_{Y;X_{1}}$, $\mathcal{U}_{Y;X_{2}}$, $\mathcal{U}_{Y;X_{3}}$), redundant ($\mathcal{R}_{Y;X}$) and synergistic ($\mathcal{S}_{Y;X}$) information rates in the time domain; the latter were obtained as the whole-band integral of the correspondent spectral functions, as well as the integral taken along two spectral bands centered around the simulated stochastic oscillations (i.e., $B_{1}=\left[0.04-0.15\right]$ Hz and $B_{2}=\left[0.15-0.4\right]$ Hz).
@@
The PIRD also favors quantification of redundancy and synergy related to the full dynamical structure of the analyzed processes, or to oscillations confined within the bands $B_{1}$ and $B_{2}$ (Fig. 5d). Redundancy arises typically from common drive (sub)structures where multiple copies of the same information are distributed, providing robustness Luppi et al. (2024a); e.g., here $X_{1}$ sends redundant information at $\sim 0.3$ Hz to both $X_{2}$ and $Y$, which is correctly detected by the significant values of $\mathcal{R}_{Y;X}$ within $B_{2}$. On the other hand, synergistic informational circuits generally emerge from common child configurations, requiring a high degree of coordination between multiple parts of the system Varley et al. (2023b); here, since $X_{1}$ and $X_{3}$ send information to $Y$ at different frequencies, synergy is detected in both bands $B_{1}$ and $B_{2}$. The balance between synergy and redundancy assessed across the full spectrum, $\Delta_{Y;X}=\mathcal{R}_{Y;X}-\mathcal{S}_{Y;X}$, indicates an an overall prevalence of redundancy within the network. Interestingly, $\Delta_{Y;X}$ indicates the presence of net synergy when assessed within $B_{1}$, due to the fact that only the source $X_{3}$ transfers information to the target $Y$ in this band, and of net redundancy when assessed within $B_{2}$, due to the common drive role of the source $X_{1}$ in this band. Therefore, the combination of common drive and common child substructures, with the source processes oscillating at different frequencies, leads to the coexistence of synergistic and redundant modes of interplay in distinct spectral bands. These complex behaviors emerging at different time scales can be detected only using frequency-specific measures of redundancy and synergy, and can be better characterized separating redundant and synergistic contributions as guaranteed by the PIRD.

### v3

@@ IV.2 Effects of changes in the network topology
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y(t_{n})}=(0.8-c){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n-1})}+(1.6-2c){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n-1})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{Y}(t_{n})}$ (35)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n})}=c{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y(t_{n-1})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{1}}(t_{n})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n})}=2c{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y(t_{n-1})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{2}}(t_{n})}$
@@
where $U_{Y}$, $U_{X_{1}}$ and $U_{X_{2}}$ are uncorrelated Gaussian white noises with zero mean and unit variance; the uncorrelation between the inputs ($\mathbf{\Sigma}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}}=\mathbf{I}$) denotes absence of instantaneous correlations. On the other hand, causal interactions are set at lag $k=1$, from $Y$ to both $X_{1}$ and $X_{2}$ with strength modulated directly by the parameter $c$, and from $X_{1}$ and $X_{2}$ to $Y$ with strength modulated inversely by the same parameter $c$. This setting allows for a progressive strengthening of the causal interactions directed from the target to the sources, and a progressive weakening of the causal interactions directed from the sources to the target, as $c$ increases in the range $[0-0.8]$. The simulation design is shown in Fig. 5a.
@@
The PIRD was performed as in the first simulation, using the approach presented in Sect. "Formulation for Gaussian processes". Remarkably, since in this simulation of processes without self-dependencies the spectral MIR functions are flat, the time- and frequency-domain PIRD are equivalent and reduce to applying the MMI criterion to the MIR. As regards the PID applied to the joint TE $I(Y(t_{n});X_{1}(t_{<n}),X_{2}(t_{<n})|Y(t_{<n}))$, it was computed exploiting the formalism linking information-theoretic measures with Granger causality (GC) measures derived from linear parametric regression models developed for joint Gaussian processes Barnett et al. (2009b); Barrett et al. (2010); Faes et al. (2017). Specifically, the MMI criterion Barrett (2015) was used to derive redundancy as the minimum information transferred from each individual source to the target; the latter was measured as half the value of the GC estimated from the VAR processes Barnett et al. (2009b), and all GC terms were computed from the VAR parameters using sub-models Faes et al. (2017). The resulting redundant TE ($\mathcal{R}$) was then used to derive the unique ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$) and synergistic ($\mathcal{S}$) amounts of information transferred from $X$ to $Y$ according to PID rules.
@@
The time domain behaviors of the PID terms are shown in Fig. 5b-f as a function of the parameters $c_{1},c_{2}$ (continuous black lines). The results evidence how the transition from the condition in which the target acts exclusively as a sink of information ($c=0$) to that in which it acts exclusively as a source of information ($c=0.8$) is thoroughly reflected by the PIRD but cannot be fully captured by the PID applied to the joint TE. Indeed, the latter cannot take time-lagged interactions directed from the target to the sources into account, and thus in the simulation where $c$ is increased it reflects only the decrease of the joint information transferred along the direction ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}\rightarrow Y$ ($J$, Fig. 5d, solid line) and the consequent drop to zero of all PID terms ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$, $\mathcal{R}$, $\mathcal{S}$, Fig. 5b,c,e,f, solid lines). On the other hand, the PIRD is applied to the MIR shared between the sources and the target, a measure which is not directional as the TE and is thus sensible to the overall information flowing among the processes of the considered dynamic system. This is reflected by values of the MIR between $Y$ and $X_{1},X_{2}$ which remain high as $c$ increases ($J$, Fig. 5d, dashed line), as also happens for the unique information rate of $X_{2}$ ($\mathcal{U}_{2}$, Fig. 5c, dashed line), and by values of redundancy rate and synergy rate shifting from the prevalence of synergy when $c=0$ to the prevalence of redundancy when $c=0.8$. The latter behavior reflects the modifications of the topological structure induced in the dynamic network, with links undergoing a transition from a common child configuration to a common drive configuration.
@@
Figure 5: Changes in the network topology have a profound impact on the multivariate information shared by multiple random processes. a) Simulation design, where $Y$ is the target process and $\{X_{1},X_{2}\}$ is the group of sources; time-lagged interactions (solid black arrows) are set varying the parameter $c\in[0-0.8]$ to simulate a transition from purely unidirectional interactions directed from $X_{1},X_{2}$ to $Y$ when $c=0$ to purely unidirectional interactions directed from $Y$ to $X_{1},X_{2}$ when $c=0.8$. (b-e) Profiles of the time domain MIR shared between the $Y$ and $\{X_{1},X_{2}\}$ and the joint TE from $\{X_{1},X_{2}\}$ to $Y$ (measuring overall interactions, here indicated as $J$) and of the unique ($\mathcal{U}_{1}$, $\mathcal{U}_{2}$), redundant ($\mathcal{R}$) and synergistic ($\mathcal{S}$) components of their decomposition measured using the PID applied to the TEs (continuous black lines) and the PIRD (dashed grey lines) as a function of the parameter $c$.
@@ IV.3 Frequency-specific coarse-grained PIRD
Figure 6: Coexistence of redundant and synergistic characters of interactions in different spectral bands elicited by frequency-specific PIRD. (a) Network structure, with $Y$ receiving from $X_{1}$, oscillating at 0.3 Hz, and $X_{3}$, oscillating at 0.1 Hz, and $X_{1}$ sending to $X_{2}$, oscillating at 0.3 Hz; all coupling coefficients are set to 1. (b) Spectral profiles of the pairwise MIR measures computed between the target $Y$ and each source $X_{i}$, $i=1,2,3$. (c,d) Time-domain values of the pairwise MIR and of the unique ($\mathcal{U}_{Y;X_{i}}$, $i=1,2,3$), redundant ($\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$) and synergistic ($\mathcal{S}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$) PIRD components integrated along the whole frequency axis ($\left[0-f_{s}/2\right]$) (left bars), and withing the bands $B_{1}=\left[0.04-0.15\right]$ Hz (middle bars), $B_{2}=\left[0.15-0.4\right]$ Hz (right bars). The balance $\Delta_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}=\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}-\mathcal{S}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ is also reported, indicating overall prevalence of redundancy and coexistence of redundancy and synergy in the two different bands $B_{1}$ and $B_{2}$.
@@
In the last simulation, we provide an illustrative example of the full-frequency and band-limited PIRD applied to a network of four nodes where different high-order behaviors emerge at different frequencies. We consider a VAR process simulated with $f_{s}=1$, where the lagged interactions between the target $Y$ and the three sources ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}=\left[X_{1},X_{2},X_{3}\right]$ are set to induce a network topology with both common drive and common child structures (Fig. 6a). The VAR process is defined as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Y(t_{n})}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n-1})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{3}(t_{n-1})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{Y}(t_{n})}$ (36)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n})}=\sum_{k=1}^{2}a_{1,k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n-k})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{1}}(t_{n})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n})}=\sum_{k=1}^{2}a_{2,k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{2}(t_{n-k})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{1}(t_{n-1})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{2}}(t_{n})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{3}(t_{n})}=\sum_{k=1}^{2}a_{3,k}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}X_{3}(t_{n-k})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{X_{3}}(t_{n})}$
@@
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{U}}=\left[U_{Y},U_{X_{1}},U_{X_{2}},U_{X_{3}}\right]$ is a vector of four zero-mean independent Gaussian white noises with unit variance. The autonomous oscillations in the three source processes $X_{1}$, $X_{2}$, $X_{3}$ are obtained placing a pair of complex-conjugate poles, with modulus $\rho$ and phase $2\pi{f}$, in the complex plane representation of each process. Here, we set $\rho=0.8,f=0.3$ Hz for the processes $X_{1}$ and $X_{2}$, so that their autonomous dynamics are determined by the coefficients $a_{1,1}=a_{2,1}=-0.494;a_{1,2}=a_{2,2}=-0.64$; similarly, we set $\rho=0.9,f=0.1$ Hz for $X_{3}$, so that its autonomous dynamics are determined by the coefficients $a_{3,1}=1.456;a_{3,2}=-0.81$. Moreover, causal interactions are set from $X_{1}$ and $X_{3}$ to $Y$ and from $X_{1}$ to $X_{2}$ at lag $k=1$, with unitary strength.
@@
The PIRD was applied computing the spectral MIRs between each individual source and the target ($i_{Y;X_{i}}(f)$, $i=1,2,3$; Fig. 6b), as well as between groups of sources and the target (e.g., $i_{Y;X_{1},X_{2}}(f)$); the spectral redundancy function was then computed as in (23), and was exploited to retrieve the unique ($\mathcal{U}_{Y;X_{1}}$, $\mathcal{U}_{Y;X_{2}}$, $\mathcal{U}_{Y;X_{3}}$), redundant ($\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$) and synergistic ($\mathcal{S}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$) information rates in the time domain; the latter were obtained as the whole-band integral of the correspondent spectral functions, as well as the integral taken along two spectral bands centered around the simulated stochastic oscillations (i.e., $B_{1}=\left[0.04-0.15\right]$ Hz and $B_{2}=\left[0.15-0.4\right]$ Hz).
@@
The PIRD also favors quantification of redundancy and synergy related to the full dynamical structure of the analyzed processes, or to oscillations confined within the bands $B_{1}$ and $B_{2}$ (Fig. 6d). Redundancy arises typically from common drive (sub)structures where multiple copies of the same information are distributed, providing robustness Luppi et al. (2024a); e.g., here $X_{1}$ sends redundant information at $\sim 0.3$ Hz to both $X_{2}$ and $Y$, which is correctly detected by the significant values of $\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ within $B_{2}$. On the other hand, synergistic informational circuits generally emerge from common child configurations, requiring a high degree of coordination between multiple parts of the system Varley et al. (2023b); here, since $X_{1}$ and $X_{3}$ send information to $Y$ at different frequencies, synergy is detected in both bands $B_{1}$ and $B_{2}$. The balance between synergy and redundancy assessed across the full spectrum, $\Delta_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}=\mathcal{R}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}-\mathcal{S}_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$, indicates an an overall prevalence of redundancy within the network. Interestingly, $\Delta_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}}$ indicates the presence of net synergy when assessed within $B_{1}$, due to the fact that only the source $X_{3}$ transfers information to the target $Y$ in this band, and of net redundancy when assessed within $B_{2}$, due to the common drive role of the source $X_{1}$ in this band. Therefore, the combination of common drive and common child substructures, with the source processes oscillating at different frequencies, leads to the coexistence of synergistic and redundant modes of interplay in distinct spectral bands. These complex behaviors emerging at different time scales can be detected only using frequency-specific measures of redundancy and synergy, and can be better characterized separating redundant and synergistic contributions as guaranteed by the PIRD.